OpenStudy (user123):

(Calculus) - I've been trying to understand this for hours. Can someone clearly explain the shell method in integration? Where does the 2 pi come from? The area of the little rectangle times 2pi equals the circumference of the surface..? And why do we just write "x"? Why is it just a random number in between 0 and the radius? Is it because it's an infinite amount of rectangles? Any help will be appreciated.

OpenStudy (tkhunny):

Volume of Right Circular Cylinder: \(\pi r^{2}\;h\) dVolume of Right Circular Cylinder: \(2\pi r\;h\;dr\) dVolume of Right Circular Cylinder: \(\pi r^{2}\;dh\)

OpenStudy (user123):

Yes, that did not satisfy my question, though.

OpenStudy (tkhunny):

Look more closely. Think hard on exactly what an integral represents and does. \(2\pi r\) is the circumference of a circle. \(h\) is the height of an open cylinder \(2\pi rh\) is the surface area of an open cylinder. Add up all the surface areas, \(\int stuff\;dr\), across all possible radii, and you have your volume. \(\pi r^{2}\) is the area of a disc. Add up all the surface areas, \(\int stuff\;dh\), across all possible heights, and you have your volume. x and y are not random. They are the limits of the radius or the height. They can't get any smaller than zero.

OpenStudy (user123):

Thanks, but I'm a bit confused because that is not the surface area of a cylinder. Well, I guess there's a difference between an open cylinder and a closed one, yes? The surface area is different..

OpenStudy (tkhunny):

I meant that is was not filled. It's just a hollow shell. Wait, isn't that why we call it he "shell method"? It IS the surface area of a cylinder - LOTS of them. More accurately, it might be called a "Representative Cylinder". Add up all the surface areas and get the volume. It's what integrals do.